Walt made an extra 9000 last year from a part time job. He invested part of the money at 9% and the rest at 8%. He made a total of $770.00 in interest. How much was invested at 8%?: Walt made an extra 9000 last year from a part time job. He invested part of the money at 9% and the rest at 8%. He made a total of $770.00 in interest. How much was invested at 8%?
Answer by ankor@dixie-net.com(3953)

Walt made an extra 9000 last year from a part time job. He invested part of the money at 9% and the rest at 8%. He made a total of $770.00 in interest. How much was invested at 8%?
:
Let x = amt invested at 8%
The total was given as $9000, therefore:
(9000-x) = amt invested at 9%
:
Write an interest equation:
:
interest at 8% + interest at 9% = $770
.08x + .09(9000-x) = 770
Multiply what's inside the brackets
.08x + 810 - .09x = 770
:
.08x - .09x = 770 - 810
:
-.01x = -40
x = (-40)/(-.01)
x = $4000 invested at 8%
:
:
Check solution ($5000 invested at 9%)
.08(4000) + .09(5000) =
320 + 450 = 770; confirms our solution